If 0.050 mol hydrochloric acid (HCl) is mixed with 0.050 mol of sodium hydroxide (NaOH) in a "coffee cup" calorimeter, the temperature of 1.41 × 102 g of the resulting solution increases from 25.2°C to 31.3°C. If the specific heat of the solution is 1.00cal/g·°C, calculate the quantity of energy involved in the reaction.
Answer #1

dT = 31.3 - 25.2 = 6.1 oC

m   = 1.41 x 10^2 g

Q = m Cp dT

Q = 1.41 x 10^2 x 4.184 x 6.1

Q = 3599 J

heat released = 3.60 kJ

enthalphy change = - Q / n

                               = -3.60 / 0.050

                             = -71.9 kJ / mole

enthalpy change = -71.9 kJ / mole